OpenStudy (anonymous):

-\frac{ 8 }{ 55}times \frac{ 3 }{ 4

OpenStudy (anonymous):

\[-\frac{ 8 }{ 55 }\times \frac{ 5 }{ 4 }\]

OpenStudy (anonymous):

You're supposed to look in the kitty corners to find out if they have any common factors

OpenStudy (anonymous):

\[ -\frac{8}{55}\times \frac{5}{4} =-\frac{2}{55}\times \frac{5}{1}=-\frac{2}{11}\times \frac{1}{1}=-\frac{2}{11}\times 1=-\frac{2}{11} \]